OpenStudy (anonymous):

Cos__=2/square root of 5

OpenStudy (lalaly):

\[\text{let the \angle }=\theta\]\[\cos^{-1} \frac{2}{\sqrt5}= \theta\]

OpenStudy (aaaa):

0.89442719